OpenStudy (anonymous):

How do you determine the domain and range of a function expressed with an equation?

OpenStudy (anonymous):

\[2-\sqrt{x}\]

ganeshie8 (ganeshie8):

do you know the domain and range of \(\sqrt{x}\) ?

OpenStudy (anonymous):

domain is \[(-\infty,\infty)\] range is \[[-3, \infty)\]

ganeshie8 (ganeshie8):

Nope. \(\sqrt{x}\) function is defined only for `nonnegative` \(x\), so the domain of \(\sqrt{x}\) is \(x \ge 0\) (assuming the given function is real valued)

ganeshie8 (ganeshie8):

In your interval notation domain of \(\sqrt{x}\) is \([0, \infty)\) and range is \([0, \infty)\)

ganeshie8 (ganeshie8):

see if you can use that info to find out the domain and range of given function

OpenStudy (anonymous):

the domain is \[[0,\infty)\] and the range is \[[2,\infty)\]???

ganeshie8 (ganeshie8):

Excellent !

ganeshie8 (ganeshie8):

how did u figre that out ? xD

OpenStudy (anonymous):

the question had a 2 in it so i figured that had something to do with either the range or the domain and basically i guessed.